Discrete Math for COMP:

(12 points) For each sequence described below, first find a recurrence relation for that sequence and then solve your recurre(c) The sequence en whose nth term is the number of edges in Kn

(12 points) For each sequence described below, first find a recurrence relation for that sequence and then solve your recurrence relation. (a) The sequence Sn where 80 = 0, si-l and, for n 〉 1, sn s the average of the previous two terms of the sequence. (b) The sequence bn whose nth term is the number of n-bit strings that don't have two zeros in a rovw
(c) The sequence en whose nth term is the number of edges in Kn
Answer #1

a) s_{n+1}=\frac{s_n+s_{n-1}}{2}

Has the char equation D^2=\frac{D}{2}+\frac{1}{2}

So D^2-\frac{D}{2}+\frac{1}{4}=\frac{3}{4}

So that \left(D-\frac{1}{2} \right )^2=\frac{3}{4}\Rightarrow D-\frac{1}{2}=\pm \frac{\sqrt{3}}{2}

And so D=\left(\frac{1\pm \sqrt{3}}{2} \right )

General solution is s_n=c_1\left(\frac{1+\sqrt{3}}{2} \right )^n+c_2\left(\frac{1-\sqrt{3}}{2} \right )^n

Using initial conditions we get s_0=c_1+c_2=0,\,s_1=c_1\left(\frac{1+\sqrt{3}}{2} \right )+c_2\left(\frac{1-\sqrt{3}}{2} \right )=1

And so c_1\left(\frac{1+\sqrt{3}}{2} \right )-c_1\left(\frac{1-\sqrt{3}}{2} \right )=1

And so c_1\left(\sqrt{3}\right )=1\Rightarrow c_1=\frac{1}{\sqrt{3}},\,c_2=-\frac{1}{\sqrt{3}}

This means s_n=\frac{1}{\sqrt{3}}\left(\frac{1+\sqrt{3}}{2} \right )^n-\frac{1}{\sqrt{3}}\left(\frac{1-\sqrt{3}}{2} \right )^n

b) b_{n+1}=b_n+b_{n-1} as the last character can be either a 0 or 1. If it is a 1, the previous n digit string can be made in b_n ways. If it is 0 the its previous character must be 1 and the previous (n-1) digit string can be made in b_{n-1} ways

Note that this is the Fibonacci recurrence with b_0=1,b_1=2 compared with F_0=0,F_1=1,F_2=1,F_3=2

So we must have b_n=F_{n+2}

c) e_1=0,e_2=1

Adding a new vertex to K_n, we need to add n new edges connecting the previous n edges to the new vertex

Thus, e_{n+1}=e_n+k so we have

e_2=e_1+2=1+2,e_3=e_2+3=(1+2)+3,\,e_k=(1+2+3\cdots )+k

And thus we have e_n=\frac{n(n+1)}{2} using the sum of natural numbers formula